### v1

# Topology of the punctual Hilbert schemes of plane curve singularities with a single Puiseux pair

Abstract
Piontkowski provided a combinatorial description of the Euler numbers and the Betti numbers of the Jacobian factors of plane curve singularities with a single Puiseux pair. In the present paper, we generalize his results to the punctual Hilbert schemes of these singularities.
Dedicated to Professor Fumio Sakai on his 70th birthday
Keywords: plane curve singularity, punctual Hilbert scheme, Euler number, Betti number
Mathematics Subject Classification (2020) 14C05, 14G10, 14H20
1 Introduction
Let $X$ be a singular plane curve over $\mathbb{C}$ with its singular point $o$. We refer to the pair $(X,o)$ as a plane curve singularity. In this paper, we consider the case where $(X,o)$ is reduced and irreducible. Let $R:=\widehat{\mathcal{O}}_{X,o}$ denote the completion of the local ring $\mathcal{O}_{X,o}$ of $X$ at $o$, and let $\Gamma$ be the semigroup of $R$. We denote by $\mathrm{Hilb}^{r}(X,o)$ the punctual Hilbert scheme of degree $r$ for the given singularity $(X,o)$. It is known that $\mathrm{Hilb}^{r}(X,o)$ admits a decomposition of the form
$\mathrm{Hilb}^{r}(X,o)=\bigsqcup_{\Delta\in\mathrm{Mod}_{r}(\Gamma)}H(\Delta)$ (1)
where $\mathrm{Mod}_{r}(\Gamma):=\{\Delta|\,\text{$\Delta$ is a $\Gamma$-semimodule with $\#(\Gamma\setminus\Delta)=r$}\}$. In general, some of $H(\Delta)$ are empty. We call the component $H(\Delta)$ the $\Delta$-subset of $\mathrm{Hilb}^{r}(X,o)$.
Piontkowski [3] studied the topology of the Jacobian factors of irreducible plane curve singularities with a single Puiseux pair. It is known that a Jacobian factor $J_{X,o}$ of an irreducible plane curve singularity $(X,o)$ is isomorphic to certain punctual Hilbert schemes (see Corollary 7 below). Therefore, $J_{X,o}$ also admits a decomposition (1). By showing that each $\Delta$-subset $H(\Delta)$ in the decomposition is isomorphic to an affine space, he described the Euler number and the Betti numbers of $J_{X,o}$. We generalize his results as follows:
Theorem 1.
Let $p$ and $q$ be coprime positive integers with $p<q$. For an irreducible plane curve singularity $(X,o)$ with $\Gamma=\langle p,q\rangle$, each $\Delta$-subset $H(\Delta)$ of $\mathrm{Hilb}^{r}(X,o)$ is isomorphic to an affine space of dimension
$\sum_{i=1}^{p-1}\#[\{(-\min\Delta+\Gamma)\cap[a_{i},a_{i}+q)\}\setminus\Delta^{(0)}]$ (2)
where $\{a_{0},\ldots,a_{p-1}\}$ is the $p$-basis of the 00-normalization $\Delta^{(0)}:=-\min\Delta+\Delta$ of $\Delta$.
Corollary 2.
Under the same assumption as in Theorem 1, the Euler number of $\mathrm{Hilb}^{r}(X,o)$ is equal to #$\#\mathrm{Mod}_{r}(\Gamma)$.
For $\Delta\in\mathrm{Mod}_{r}(\Gamma)$, define the codimension of $H(\Delta)$ as $\mathrm{codim}\,H(\Delta):=\dim\mathrm{Hilb}^{r}(X,o)-\dim H(\Delta).$ We also set
$\mathcal{H}_{r,d}:$ =$=\{H(\Delta)|\,\Delta\in\mathrm{Mod}_{r}(\Gamma)\text{ and }\dim H(\Delta)=d\},$
$\mathcal{H}^{d}_{r}:$ =$=\{H(\Delta)|\,\Delta\in\mathrm{Mod}_{r}(\Gamma)\text{ and }\mathrm{codim}\,H(\Delta)=d\}.$
Theorem 3.
Let $(X,o)$ be the same as in Theorem 1. The odd ((((co-)))) homology groups of $\mathrm{Hilb}^{r}(X,o)$ vanish. The even ((((co-)))) homology groups of $\mathrm{Hilb}^{r}(X,o)$ are free abelian groups with Betti numbers
$h_{2d}(\mathrm{Hilb}^{r}(X,o))=\#\mathcal{H}_{r,d}$ and $h^{2d}(\mathrm{Hilb}^{r}(X,o))=\#\mathcal{H}^{d}_{r}$.
Remark 4.
In [3], Piontkowski originally proved Theorem 1, Corollary 2, and Theorem 3 for the case where $r\geq c$.
This paper is structured as follows: Sections 2, 3, and 4 summarize the foundational concepts required for the proofs. Section 2 recalls the properties of punctual Hilbert schemes of curve singularities. Section 3 introduces definitions and facts related to $\Gamma$-semimodules. Section 4 discusses computational techniques involving Gröbner bases. In Section 5, we prove the main theorems. Section 6 presents examples, and Section 7 provides remarks on related results.
2 Punctual Hilbert schemes of curve singularities
In this section, we recall properties of punctual Hilbert schemes established in [4]. Although we focus on irreducible plane curve singularities here, the notions presented here hold in more general situations.
Let $(X,o)$ be an irreducible plane curve singularity. By Puiseux’s theorem, there exist coprime positive integers $p$ and $q$ such that $R=\mathbb{C}[[t^{p},\phi]]$ where $\phi=t^{q}+\text{higher order terms}$ and $p<q$. The normalization $\overline{R}$ of $R$ is isomorphic to $\mathbb{C}[[t]]$. Let $\nu$ be the natural valuation $\nu:\overline{R}\setminus\{0\}\rightarrow\mathbb{Z}_{\geq 0}$ defined by $\nu(f)=\mathrm{ord}_{t}(f)$, and set $\nu(0)=\infty$. We call $\Gamma:=\nu(R)$ the semigroup of $R$. A semigroup $\Gamma$ generated by positive integers $\gamma_{1},\gamma_{2},\ldots,\gamma_{l}$ is denoted as $\Gamma=\langle\gamma_{1},\gamma_{2},\ldots,\gamma_{l}\rangle$. The $\delta$-invariant of $R$ is defined as $\delta:=\dim_{\mathbb{C}}(\overline{R}/R)$. The conductor $c$ of $\Gamma$ is the smallest integer in $\Gamma$ such that $c-1\notin\Gamma$ and $c+n\in\Gamma$ for any $n\in\mathbb{N}$. It satisfies $\delta+1\leq c\leq 2\delta$ and $c=2\delta$ if and only if $R$ is Gorenstein (see [6]).
For a positive integer $a$, we denote by $(t^{a})$ an ideal of $\overline{R}$. Let $\mathrm{Gr}\left(\delta,\overline{R}/(t^{2\delta})\right)$ be the Grassmannian which consists of $\delta$-dimensional linear subspaces of $\overline{R}/(t^{2\delta})$. For $M\in\mathrm{Gr}\left(\delta,\overline{R}/(t^{2\delta})\right)$, we define a multiplication by $R\times M\ni(f,m+(t^{2\delta}))\mapsto fm+(t^{2\delta})\in M$. The set
$J_{X,o}:=\bigg{\{}M\in\mathrm{Gr}\left(\delta,\overline{R}/(t^{2\delta})\right)\bigg{|}\,\begin{array}[]{l}\text{$M$ is an $R$-submodule with respect to}\\ \text{ the above multiplication.}\end{array}\bigg{\}}$
is called the Jacobian factor of $(X,o)$, which was introduced by Rego in [5]. For any non-negative integer $r$, set
$\mathcal{I}_{r}:=\{I\subset R|\,\text{$I$ is an ideal of $R$ with $\dim R/I=r$}\}.$
Observe that any element $I$ in $\mathcal{I}_{r}$ satisfies $(t^{r+2\delta})\cap R\subseteq I\subseteq(t^{r})\cap R$. It follows that $(t^{2\delta})\subseteq t^{-r}I\subseteq\overline{R}$ and $\dim_{k}(\overline{R}/t^{-r}I)=\dim_{k}(\overline{R}/I)-r=\delta$. Using these properties, Pfister and Steenbrink defined a map
$\varphi_{r}:\mathcal{I}_{r}\rightarrow J_{X,o}\subset\mathrm{Gr}(\delta,\overline{R}/(t^{2\delta}))$ (3)
by $\varphi_{r}(I)=t^{-r}I/(t^{2\delta})$ (see [4]). We call $\varphi_{r}$ the $\delta$-normalized embedding. It has the following properties:
Proposition 5 ([4], Theorem 3).
The $\delta$-normalized embedding $\varphi_{r}$ is injective for any non-negative integer $r$. It is also a bijection for $r\geq c$. The punctual Hilbert scheme $\mathrm{Hilb}^{r}(X,o)$ is Zariski closed in $J_{X,o}$.
Definition 6.
We call $\mathrm{Hilb}^{r}(X,o):=\varphi_{r}(\mathcal{I}_{r})$ the punctual Hilbert scheme of degree $r$ for $(X,o)$.
The following fact follows from Proposition 5:
Corollary 7.
For any positive integer $r$ with $r\geq c$, we have
$\mathrm{Hilb}^{r}(X,o)\cong\mathrm{Hilb}^{c}(X,o)\cong J_{X,o}.$
3 $\Gamma$-semimodules
Let $\Gamma$ be a semigroup. A subset $\Delta\subset\mathbb{Z}$ is called a $\Gamma$-semimodule if it satisfies $\Delta+\Gamma\subseteq\Delta$. It is straitforward to see that, for any ideal $I$ of $R$, the set $\Gamma(I):=\{\nu(f)|\,f\in I\}$ is a $\Gamma$-semimodule. If integers $a_{1},\ldots,a_{s}$ generate $\Delta$ (i.e., $\Delta=\sum_{i=1}^{s}(a_{i}+\Gamma)$ holds), we write $\Gamma=\langle a_{1},\ldots,a_{s}\rangle_{\Gamma}$. Two $\Gamma$-semimodules $\Delta_{1}$ and $\Delta_{2}$ are isomorphic if there exists an integer $a$ such that $\Delta_{1}=a+\Delta_{2}=\{a+d|\,d\in\Delta_{2}\}$. We define two special normalizations of a $\Gamma$-semimodule $\Delta$: We define the 00-normalization of $\Delta$ as $\Delta^{(0)}:=-\min\Delta+\Delta$. In general, $\Delta$ is called 00-normalized if $\min{\Delta}=0$. On the other hand, $\Delta$ is said to be $\delta$-normalized if #$\#(\mathbb{N}\setminus\Delta)=\delta$. For $I\in\mathcal{I}_{r}$, there exists an element $\Delta$ of $\mathrm{Mod}_{r}(\Gamma)$ such that $\Delta=\Gamma(I)$. From the inclusion $(t^{2\delta})\subset t^{-r}I$ and the definition of $\varphi_{r}$ in (3), it follows that $-r+\Delta$ is $\delta$-normalized, and #$\#\{(-r+\Delta)\cap[0,2\delta-1]\}=\delta$ holds. We call $-r+\Delta$ the $\delta$-normalization of $\Delta$ and denote it by $\Delta^{(\delta)}$.
In [3], Piontkowski introduced special generators of a $\Gamma$-semimodule.
Definition 8 ([3]).
Let $\Gamma=\langle p,q\rangle$ where $p<q$ and $\gcd(p,q)=1$. The $p$-basis of a $\Gamma$-semimodule $\Delta$ is the unique set $\{a_{0},a_{1},\ldots,a_{p-1}\}$ satisfying
$\Delta=\bigcup_{i=0}^{p-1}(a_{i}+p\mathbb{N})\text{\ \ and\ \ }a_{i}\equiv iq\text{ $(\mathrm{mod}\ p)$}.$
In particular, we have $\Delta=\langle a_{0},\ldots,a_{p-1}\rangle_{\Gamma}$.
Throughout this paper, we consider the $p$-basis only for 00-normalized $\Gamma$-semimodules and therefore we always set $a_{0}=0$. Using the inclusion $\mathbb{N}q\subset\Gamma\subset\Delta^{(0)}$, we find non-negative integers $\alpha_{1},\ldots,\alpha_{p-1}\in\mathbb{N}$ such that
$a_{0}=0,\,a_{1}=q-\alpha_{1}p,\,a_{2}=2q-\alpha_{2}p,\,\ldots,\,a_{p-1}=(p-1)q-\alpha_{p-1}p.$ (4)
We also set $\alpha_{0}=0$. These satisfy $0\leq\alpha_{1}\leq\alpha_{2}\leq\cdots\leq\alpha_{p-1}<q.$
The set $\mathcal{I}_{r}$ can be decomposed in terms of $\Gamma$-semimodules.
Lemma 9 ([7], Proposition 6).
We have
$\mathcal{I}_{r}=\bigsqcup_{\Delta\in\mathrm{Mod}_{r}(\Gamma)}\mathcal{I}(\Delta)$ (5)
where $\mathcal{I}(\Delta):=\{I\in\mathcal{I}_{r}|\,\Gamma(I)=\Delta\}$.
By setting $H(\Delta):=\varphi_{r}(\mathcal{I}(\Delta))$, the stratification (1) of $\mathrm{Hilb}^{r}(X,o)$ follows from Lemma 9.
Remark 10.
[…]
4 Gröbner bases
We begin by recalling some key facts about Gröbner bases, primarily following [2]. Let $R=\mathbb{C}[[x_{1}(t),\ldots,x_{l}(t)]]$ be a subring of $\mathbb{C}[[t]]$ such that $\dim_{\mathbb{C}}\mathbb{C}[[t]]/R<\infty$, and let $M\subset\mathbb{C}[[t]]$ be an $R$-module. As in Section 2, let $\nu$ be the natural valuation $\nu:\mathbb{C}[[t]]\setminus\{0\}\rightarrow\mathbb{Z}_{\geq 0}$. We also set $\nu(0)=\infty$. Put $\Gamma:=\nu(R)$ and $\Gamma(M):=\nu(M)$. Clearly, $\Gamma(M)$ is a $\Gamma$-semimodule. We consider the local order $\nu(1)\succ\nu(t)\succ\nu(t^{2})\succ\cdots$. For $f\in M$, we denote by ${\small\mathrm{LC}}(f)$ (resp. ${\small\mathrm{LT}}(f))$ the leading coefficient of $f$ (resp. the leading term of $f$) with respect to this order.
Definition 11.
A subset $G=\{g_{1},\ldots,g_{m}\}$ of $R$ is called a SAGBI basis ((((Subalgebra Analog to Gröbner Bases for Ideals)))) if, for any $f\in R$, there exists a multi-index $(\beta_{1},\ldots,\beta_{m})\in\mathbb{Z}^{m}_{\geq 0}$ such that $\mathrm{\small LT}(f)=\mathrm{\small LT}(g_{1}^{\beta_{1}}\cdots g_{m}^{\beta_{m}})$.
Definition 11 implies the following:
Theorem 12.
A SAGBI basis $G$ of R generates $R$ as $\mathbb{C}$-algebra.
Definition 13.
Let $G$ be a SAGBI basis for $R$ and let $H$ be a subset of $M$. The pair $(G,H)$ is called a standard basis of $M$ if, for any $f\in M$, there exist $h\in H$ and a multi-index $(\beta_{1},\ldots,\beta_{m})\in\mathbb{Z}^{m}_{\geq 0}$ such that $\mathrm{\small LT}(f)=\mathrm{\small LT}(g_{1}^{\beta_{1}}\cdots g_{m}^{\beta_{m}}h)$.
Similar to the above, the following theorem holds:
Theorem 14.
Let $(G,H)$ be a standard basis of $M$. The set $H$ generates $M$ as $R$-module.
The following proposition follows from Definition 13.
Proposition 15.
Let $G=\{g_{1},\ldots,g_{m}\}\subset R$ be a SAGBI basis for $R$ and let $H=\{h_{1},\ldots,h_{n}\}$ be a subset of $M$. Then $(G,H)$ is a standard basis of $M$ if and only if $\Gamma=\langle\nu(g_{1}),\ldots,\nu(g_{m})\rangle$ and $\Gamma(M)=\langle\nu(h_{1}),\ldots,\nu(h_{n})\rangle_{\Gamma}$.
Theorem 16.
Let $G$ and $H$ be as in Proposition 15. If $(G,H)$ is a standard basis of $M$, then any $f\in\mathbb{C}[[t]]$ can be written as
$f=q_{1}h_{1}+\cdots+q_{n}h_{n}+r$ (6)
where $q_{1},\ldots,q_{n}\in\mathbb{C}[[g_{1},\ldots,g_{m}]]$ and $r=\sum_{j\notin\Gamma(M)}c_{j}t^{j}.$
We call $r$ in (6) the reduction of $f$ modulo $(G,H)$, and denote it by $R(f,G,H)$.
Proof.
We claim that the following algorithm yields (6):
Division Algorithm
Input: $f\in R$, $G=\{g_{1},\ldots,g_{m}\}$, $H=\{h_{1},\ldots,h_{n}\}$
Output: $q_{1},\ldots,q_{n}$, $r$
Define: $q_{1}:=0,\ldots,q_{n}:=0$, $r=0$, $p:=f$
WHILE $p\neq 0$ DO
 If $\nu(p)\in\Gamma(M)$, THEN find smallest $l$ such that
$\mathrm{\small LT}(p)=\mathrm{\small LT}(g_{1}^{\alpha_{1}}\cdots g_{m}^{\beta_{m}}h_{l})\text{ for some }(\beta_{1},\ldots,\beta_{m})\in\mathbb{Z}^{m}_{\geq 0}$
  $q_{l}:=q_{l}+g_{1}^{\beta_{1}}\cdots g_{m}^{\beta_{m}}$
  $p:=p-g_{1}^{\beta_{1}}\cdots g_{m}^{\beta_{m}}h_{l}$
  $r:=r$
ELSE
  $p:=p-\mathrm{\small LT}(p)$
  $r:=r+\mathrm{\small LT}(f)$
We allow the algorithm to proceed infinitely many steps. Since the order $\nu(p)$ strictly increases at every step, $p$ converges to 0. It is also obvious that the orders of all terms in the final reduction $r$ are not in $\Gamma(M)$. For further details, refer to the proof of Theorem 3 in Chapter 2, Section 3 of [1]. ∎
The reduction $R(f,G,H)$ has the following property:
Lemma 17.
Let $(G,H)$ be a standard basis of $M$. For any $f\in M$, the reduction $R(f,G,H)$ is unique no matter how the elements of $G$ and $H$ are listed.
Proof.
The proof of this lemma is similar to that of Proposition 1 in Chapter 2, Section 6 of [1]. So we omit it. ∎
For $f_{1},f_{2}\in M$, there exist multi-indices $(\beta_{1},\ldots,\beta_{m}),(\gamma_{1},\ldots,\gamma_{m})\in\mathbb{Z}^{m}_{\geq 0}$ and elements $h_{1},h_{2}\in H$ such that $\mathrm{\small LT}(f_{1})=\mathrm{\small LT}(g_{1}^{\beta_{1}}\cdots g_{m}^{\beta_{m}}h_{1})$ and $\mathrm{\small LT}(f_{2})=\mathrm{\small LT}(g_{1}^{\gamma_{1}}\cdots g_{m}^{\gamma_{m}}h_{2})$. The $S$-process for $f_{1},f_{2}\in M$ is defined as
$S(f_{1},f_{2}):=g_{1}^{\gamma_{1}}\cdots g_{m}^{\gamma_{m}}h_{2}f_{1}-g_{1}^{\beta_{1}}\cdots g_{m}^{\beta_{m}}h_{1}f_{2}.$
Among all such expressions, we define the one with minimal order as the minimal $S$-process of $f_{1}$ and $f_{2}$, denoted by $S_{\mathrm{min}}(f_{1},f_{2})$.
Proposition 18 ([2], Theorem 2.3).
Let $G=\{g_{1},\ldots,g_{m}\}\subset R$ be a SAGBI basis for $R$ and let $H=\{h_{1},\ldots,h_{n}\}$ be a subset of $M$. The pair $(G,H)$ is a standard basis of $M$ if and only if $R(S_{\mathrm{min}}(h_{i},h_{j}),G,H)=0$ for all $h_{i},h_{j}\in H$.
5 Proofs of the Main Results
As noted in Remark 4, Piontkowski [3] originally proved Theorem 1, Corollary 2, and Theorem 3 in the case $r\geq c$. The following result about syzygies was used in his proof:
Lemma 19 ([3], Proposition 5).
Let $\Gamma=\langle p,q\rangle$, and let $\Delta=\cup^{p-1}_{i=0}(a_{i}+p\mathbb{N})$ be a $\Gamma$-semimoudlue with $p$-basis $\{a_{0},\ldots,a_{p-1}\}$. Consider a graded algebra $\mathbb{C}[\Gamma]:=\bigoplus_{\gamma\in\Gamma}\mathbb{C}t^{\gamma}$, and a graded $\mathbb{C}[\Gamma]$-algebra $\mathbb{C}[\Delta]:=\bigoplus_{\gamma\in\Delta}\mathbb{C}t^{\gamma}$ generated by $(t^{a_{0}},\ldots,t^{a_{p-1}})$. Then the syzygies of this $p$-tuple are minimally generated by the following vectors::::
$v_{0}:$ $=(t^{q},-t^{\alpha_{1}p},0,\ldots,0),$
$v_{1}:$ $=(0,t^{q},-t^{(\alpha_{2}-\alpha_{1})p},0,\ldots,0),$
[…]
$v_{p-2}:$ $=(0,\ldots,0,t^{q},-t^{(\alpha_{p-1}-\alpha_{p-2})p}),$
$v_{p-1}:$ $=(-t^{(q-\alpha_{p-1})p},0,\ldots,0,t^{q})$
[…]
Before proving the main theorems, we briefly summarize Piontkowski’s approach.
Outline of Piontkowski’s proofs. If $r\geq c$, then we have $\mathrm{Hilb}^{r}(X,o)\cong\mathrm{Hilb}^{c}(X,o)\cong J_{X,o}$ by Corollary 7. Recall that points of $\mathrm{Hilb}^{c}(X,o)$ correspond bijectively to elements $I\in\mathcal{I}_{c}$ via $\varphi_{c}$. Moreover, each $I\in\mathcal{I}_{c}$ corresponds to $t^{-d}I$ where $d:=\min I$. If $\Delta:=\nu(I)$, then $\nu(t^{-d}I)=\Delta^{(0)}$. Thus, Piontkowski identified $t^{-d}I$ with a point in the $\Delta$-subset $H(\Delta)$ of $J_{X,o}=\mathrm{Hilb}^{c}(X,o)$. In his proof, the following generators of $t^{-d}I$ were used:
$h_{i}$ $=t^{a_{i}}+\sum_{\tiny a_{i}+k\notin\Delta^{(0)},\ k\geq 0}\lambda_{i,k}t^{a_{i}+k},\ (i=0,\ldots,p-1)$ (7)
where $\{a_{0},a_{1},\ldots,a_{p-1}\}$ is the $p$-basis of $\Delta^{(0)}$. He verified the conditions for $\Delta^{(0)}=\langle\nu(h_{0}),\ldots,\nu(h_{p-1})\rangle_{\Gamma}$. Set $G=\{t^{p},\phi\}$ where $\phi=t^{q}+\text{hight order terms}$, and $H=\{h_{0},\ldots,h_{p-1}\}$. Then $\Delta^{(0)}=\langle\nu(h_{0}),\ldots,\nu(h_{p-1})\rangle_{\Gamma}$ holds if and only if $(G,H)$ is a standard basis of $R$-module $t^{-d}I$ by Proposition 15. Moreover, by Proposition 18 and Lemma 19, $(G,H)$ is a standard basis of $t^{-d}I$ if and only if
$R(S_{\mathrm{min}}(h_{i},h_{i+1}),G,H)=0\text{ for }i=0,\ldots,p-2,$ (8)
$R(S_{\mathrm{min}}(h_{p-1},h_{0}),G,H)=0$ (9)
hold. The minimal $S$-processes are given by
$S_{\mathrm{min}}(h_{i},h_{i+1})$ $=\phi h_{i}-t^{(\alpha_{i+1}-\alpha_{i})p}h_{i+1}\text{ for }i=0,\ldots,p-2,$
$S_{\mathrm{min}}(h_{p-1},h_{0})$ $=\phi h_{p-1}-t^{(q-\alpha_{p-1})p}h_{0}.$
Using (8) and (9), he analyzed the coefficients of $S$-processes and concluded that the number of independent coefficients in $h_{i}$’s such that $\Delta^{(0)}=\langle\nu(h_{0}),\ldots,\nu(h_{p-1})\rangle_{\Gamma}$ is given by
$\sum_{i=0}^{p-1}\#\{[a_{i},a_{i}+q]\setminus\Delta^{(0)}\}.$ (10)
Hence, each $\Delta$-subset $H(\Delta)$ of $J_{X,o}=\mathrm{Hilb}^{c}(X,o)$ is an affine cell of dimension (10). When $r\geq c$, we see that $\min\Delta\geq c$, and so $(-\min\Delta+\Gamma)\cap[a_{i},a_{i}+q]=[a_{i},a_{i}+q]$, making (2) equal to (10).
Once the decomposition (1) is shown to be an affine cell decomposition, it follows that Euler number of $J_{X,o}$ equals the number of $\Delta$-subsets in (1). Moreover, these $\Delta$-subsets form a CW complex. So the Betti numbers follow from standard (co-)homology theory (see [3] for details). $\square$
Proofs of Theorem 1, Corollary 2 and Theorem 3. Piontkowski’s method applies directly to our general case without modification. We note one remark about the difference between the dimensions (2) and (10) of $\Delta$-subsets. Let $I\in I(\Delta)\subset\mathcal{I}_{r}$, and set $d=\min\{I\}$. Consider the generators $t^{d}h_{i}$ $(i=0,\ldots,p-1)$ of $I$, which can be written as
$t^{d}h_{i}=t^{b_{i}}+\sum_{\tiny b_{i}+k\notin\Delta,\ k\geq 0}\lambda_{i,k}t^{b_{i}+k}$ (11)
where $b_{i}:=a_{i}+d$ and $(i=0,\ldots,p-1)$. Every ideal in $I(\Delta)$ has generators of the same form as (11). The numbers of coefficients in (7) and (11) are same, since the sets $\mathbb{N}\setminus\Delta^{(0)}$ and $\{n\in\mathbb{N}\setminus\Delta|\,n>d\}$ are in one-to-one correspondence. Piontokowski’s idea was to count the number of independent coefficients in (7) (equivalently, in (11)). This count gives the dimension of $H(\Delta)$ as an affine space. If $r<c$, then (11) may contain some terms whose orders are not in $\Gamma$. We claim that the coefficients of such terms are not independent. Setting $A:=\{\beta_{i}\notin\Gamma|\,\beta_{i}>q\}:=\{\beta_{1},\ldots,\beta_{s}\}$, we may assume that $\phi$ is of form $t^{q}+\sum_{\beta_{i}\in A}c_{i}t^{\beta_{i}}$. The set $\{t^{p},\phi\}$ is a standard basis of $R$ in ordinary sense. Reducing $t^{d}h_{i}$ by $\{t^{p},\phi\}$, the final reduction of $t^{d}h_{i}$ must be 0, since $t^{d}h_{i}\in R$. This implies that, for each term $\lambda_{i,k}t^{b_{i}+k}$ with $b_{i}+k\notin\Gamma$, there exists polynomial $f_{i,k}(x_{1},\ldots,x_{s})$ in $\mathbb{C}[x_{1},\ldots,x_{s}]$ such that $\lambda_{i,k}=f_{i,k}(c_{1},\ldots,c_{s})$ (i.e. the coefficient $\lambda_{i,k}$ is determined by $\varphi$, and hence are not independent). This fact explains the difference between (2) and (10). $\square$
6 Examples
The punctual Hilbert schemes of curve singularity of type $A_{2l}$ ((((i.e., the curve singularity with $R=\mathbb{C}[[t^{2},t^{2l+1}]])$ was studied in [8]. For this singularity, we have $\Gamma=\langle 2,2l+1\rangle$, $\delta=l$, and $c=2l$. Let $e(\mathrm{Hilb}^{r}(X,o))$ denote the Euler number of $\mathrm{Hilb}^{r}(X,o)$. Let [ $\cdot$ ] denote the greatest integer function, where for a real number $a$, the value [$a$] is the largest integer satisfying $[a]\leq a$. The following example is obtained from the results in [8]:
Example 20.
Let $(X,o)$ be the $A_{2l}$-singularity. The Euler numbers of the punctual Hilbert schemes $\mathrm{Hilb}^{r}(X,o)$ are given in the following table::::
$r$ $10\leq r\leq 2l-$ 1 $r\geq 2l$
$e(\mathrm{Hilb}^{r}(X,o))$ $[r/2]+1$ $l+1$
The Betti numbers of $\mathrm{Hilb}^{r}(X,o)$ are as follows::::
$h_{2d}(\mathrm{Hilb}^{r}(X,o))=h^{2d}(\mathrm{Hilb}^{r}(X,o))=1\quad(d\in\{0,1,\ldots,[r/2]\})$
The punctual Hilbert schemes of curve singularities of type $E_{6}$ and $E_{8}$ ((((i.e. the curve singularities with $R=\mathbb{C}[[t^{3},t^{4}]]$ and $R=\mathbb{C}[[t^{3},t^{5}]]$ respectively) were studied in [7]. For the $E_{6}$-singularity, the fundamental invariants are $\Gamma=\langle 3,4\rangle$, $\delta=3$ and $c=6$. On the other hand, we have $\Gamma=\langle 3,5\rangle$, $\delta=4$ and $c=8$. The following examples are derived from the results in [7]:
Example 21.
Let $(X,o)$ be the $E_{6}$-singularity. The Euler numbers of the punctual Hilbert schemes $\mathrm{Hilb}^{r}(X,o)$ are given in the following table::::
$r$ 00 $1$ $2$ $3$ $4$ $5$ $r\geq 6$
$e(\mathrm{Hilb}^{r}(X,o))$ $1$ $1$ $2$ $3$ $4$ $4$ $5$
The Betti numbers of $\mathrm{Hilb}^{r}(X,o)$ are also given in the following tables::::
$r$ $h_{0}$ $h_{2}$ $h_{4}$ $h_{6}$
00 $1$
$1$ $1$
$2$ $1$ $1$
$3$ $1$ $1$ 1
$4$ $1$ $1$ 2
$5$ $1$ $1$ 2
$6$ $1$ $1$ 2 1
$r$ $h^{0}$ $h^{2}$ $h^{4}$ $h^{6}$
00 $1$
$1$ $1$
$2$ $1$ $1$
$3$ $1$ $1$ 1
$4$ $2$ $1$ 1
$5$ $2$ $1$ 1
$6$ $1$ $2$ 1 1
Example 22.
Let $(X,o)$ be the $E_{8}$-singularity. The Euler numbers of the punctual Hilbert schemes $\mathrm{Hilb}^{r}(X,o)$ are given in the following table::::
$r$ 00 $1$ $2$ $3$ $4$ $5$ $6$ $7$ $r\geq 8$
$e(\mathrm{Hilb}^{r}(X,o))$ $1$ $1$ $2$ $3$ $4$ $5$ $6$ $6$ $7$
The Betti numbers of $\mathrm{Hilb}^{r}(X,o)$ are also given in the following tables::::
$r$ $h_{0}$ $h_{2}$ $h_{4}$ $h_{6}$ $h_{8}$
00 $1$
$1$ $1$
$2$ $1$ $1$
$3$ $1$ $1$ 1
$4$ $1$ $1$ 2
$5$ $1$ $1$ 2 1
$6$ $1$ $2$ 1 2
$7$ $1$ $1$ $2$ $2$
$8$ $1$ $1$ $2$ $2$ $1$
$r$ $h^{0}$ $h^{2}$ $h^{4}$ $h^{6}$ $h^{8}$
00 $1$
$1$ $1$
$2$ $1$ $1$
$3$ $1$ $1$ 1
$4$ $2$ $1$ 1
$5$ $1$ $2$ 1 1
$6$ $2$ $1$ 2 1
$7$ $2$ $2$ 1 1
$8$ $1$ $2$ 2 1 1
7 Remarks on the results similar to Theorem 1
In this section, we discuss the results of Pfister and Steenbrink in [4]. Here we consider a monomial curve singularity $(X,o)$.
Definition 23 ([4], Definition 8).
We call a curve singularity $(X,o)$ with $R=\mathbb{C}[[t^{a_{1}},\cdots,t^{a_{n}}]]$ for some positive integers $a_{1},\ldots,a_{n}$ a monomial curve singularity.
Without loss of generality, we assume that $\gcd(a_{1},\ldots,a_{n})=1$ in Definition 23. Additionally, we focus on special semigroups defined as follows:
Definition 24 ([4], Definition 9).
A semigroup $\Gamma$ is called monomial if $0\in\Gamma$, #$\#(\mathbb{N}\setminus\Gamma)<\infty$, and any reduced, irreducible curve singularity with $\Gamma$ is a monomial curve singularity.
The monomial semigroups were completely determined in [4].
Theorem 25 ([4], Theorem 10).
A monomial semigroup is monomial if and only if it is one of the following three types::::
$\{im|\,i=0,1,\ldots,s\}\cup[sm+b,\infty)\text{ with }1\leq b<m,\ s\geq 1,$
$\{0\}\cup[m,m+r-1]\cup[m+r+1,\infty)\text{ with }2\leq r\leq m-1,$
$\{0,m\}\cup[m+2,2m]\cup[2m+2,\infty)\text{ with }m\geq 3$
Example 26.
Semigroups $\langle 2,2l+1\rangle$ with $l\geq 1$, $\langle 3,4\rangle$, and $\langle 3,5\rangle$ are all monomial semigroups that can be realized as plane curve singularities.
For a $\Gamma$-semimodule $\Delta$, let $S$ be the set of minimal generators of $\Delta^{(\delta)}$. Set $S^{\prime}:=S\cap[0,2\delta-1]$. For each $\gamma\in S^{\prime}$, define $J_{\gamma}:=[\gamma+1,2\delta-1]\setminus\Delta^{(\delta)}$.
Theorem 27 ([4], Theorem 11).
Let $(X,o)$ be a curve singularity with a monomial semigroup. For $I\in\mathcal{I}_{r}$, there exist uniquely determined $u_{\gamma,j}\in\mathbb{C}$ such that, as $R$-submodule, $\varphi_{r}(I)$ is generated by
$g_{\gamma}=t^{\gamma}+\sum_{j\in J_{\gamma}}u_{\gamma,j}t^{j},(\gamma\in S^{\prime}).$
Corollary 28 ([4], Corollary of Theorem 11).
For a curve singularity $(X,o)$ with a monomial semigroup, a $\Delta$-subset $H(\Delta)$ of $\mathrm{Hilb}^{r}(X,o)$ is an affine space of dimension $\sum_{\gamma\in S^{\prime}}\#J_{\gamma}$.
Remark 29.
Theorem 1 is stated in terms of $\Delta^{(0)}$, whereas Theorem 27 and Corollary 28 are expressed using $\Delta^{(\delta)}$.
However, Theorem 27 and Corollary 28 do not hold in general. Suppose that Theorem 27 and Corollary 28 do hold for the $E_{6}$-singularity. Recall that $R=\mathbb{C}[[t^{3},t^{4}]]$, $\Gamma=\langle 3,4\rangle$ and $\delta=3$ for this singularity. Consider a $\Gamma$-semimodule $\Delta=\langle 4,6,7\rangle_{\Gamma}$. It is easy to verify that $\Delta\in\mathrm{Mod}_{2}(\Gamma)$. So its $\delta$-normarization is give by $\Delta^{(\delta)}=-2+\Delta=\langle 2,4,5\rangle_{\Gamma}$. Let $I$ be an element of $\mathcal{I}(\Delta)$. It follows from Theorem 27 that $\varphi_{2}(I)$ is generated by $g_{2}=t^{2}+u_{2,3}t^{3}$, $g_{4}=t^{4}$ and $g_{5}=t^{5}$ as $R$-submodule. By Corollary 28, we also have $\dim H(\Delta)=1$. However, we see that $\mathcal{I}(\Delta)=\{(t^{4},t^{6},t^{7})\}$, which means $H(\Delta)$ consist of a single point. Therefore, $\dim H(\Delta)=0$. Moreover, it is obvious that $\varphi_{2}((t^{4},t^{6},t^{7}))$ is generated by $t^{2}$, $t^{4}$, and $t^{5}$. These contradict the assumption.